# RCAEval: A Benchmark for Root Cause Analysis of Microservice Systems with Telemetry Data

### v4

@@ Abstract.
Root cause analysis (RCA) for microservice systems has gained significant attention in recent years. However, there is still no standard benchmark that includes large-scale datasets and supports comprehensive evaluation environments. In this paper, we introduce RCAEval, an open-source benchmark that provides datasets and an evaluation environment for RCA in microservice systems. First, we introduce three comprehensive datasets comprising 735 failure cases collected from three microservice systems, covering various fault types observed in real-world failures. Second, we present a comprehensive evaluation framework that includes fifteen reproducible baselines covering a wide range of RCA approaches, with the ability to evaluate both coarse-grained and fine-grained RCA. RCAEval is designed to support both researchers and practitioners. We hope that this ready-to-use benchmark will enable researchers and practitioners to conduct extensive analysis and pave the way for robust new solutions for RCA of microservice systems.
@@ 1. Introduction
In this work, we introduce RCAEval, a comprehensive benchmark, consisting of three RCA datasets and a comprehensive evaluation environment with fifteen baselines. First, our datasets include 735 failure cases collected from three microservice systems, covering 11 fault types observed in real-world failures. We collected multi-source telemetry data (i.e., metrics, logs, and traces), supporting a variety of RCA approaches (e.g., metric-based, trace-based, multi-source RCA). Second, we release our evaluation framework as an open-source library, which includes reproducible RCA baselines for benchmarking. The prior version of this framework was used to evaluate metric-based RCA (et al., 2024). In this work, we have upgraded it to support trace-based and multi-source RCA. We also provide preliminary experiments, highlighting the need for further investigation on a robust RCA approach. We have open-sourced our benchmark, including the datasets and the evaluation framework at https://github.com/phamquiluan/RCAEval.

### v5

@@ Abstract.
Root cause analysis (RCA) for microservice systems has gained significant attention in recent years. However, there is still no standard benchmark that includes large-scale datasets and supports comprehensive evaluation environments. In this paper, we introduce RCAEval, an open-source benchmark that provides datasets and an evaluation environment for RCA in microservice systems. First, we introduce three comprehensive datasets comprising 735 failure cases collected from three microservice systems, covering various fault types observed in real-world failures. Second, we present a comprehensive evaluation framework that includes fifteen reproducible baselines covering a wide range of RCA approaches, with the ability to evaluate both coarse-grained and fine-grained RCA. We hope that this ready-to-use benchmark will enable researchers and practitioners to conduct extensive analysis and pave the way for robust new solutions for RCA of microservice systems.
@@ 1. Introduction
In this work, we introduce RCAEval, a benchmark including three datasets and a comprehensive evaluation environment. First, our datasets include 735 failure cases collected from three systems, covering 11 fault types observed in real-world failures. We collected multi-source telemetry data (i.e., metrics, logs, and traces), supporting a variety of RCA approaches (e.g., metric-based, multi-source RCA). Second, we open-source our evaluation framework, which includes several reproducible baselines. The prior version of this framework was used to evaluate metric-based RCA (et al., 2024). In this work, we have upgraded it to support trace-based and multi-source RCA. We also provide preliminary experiments, highlighting the need for further investigation on a robust RCA approach. We have open-sourced our benchmark, including the datasets and the evaluation framework at https://github.com/phamquiluan/RCAEval.